# Forgetting is Everywhere

### v1

@@ 3.2 Induced future distributions
The construction above motivates a predictive perspective of learning, conceptually inspired by the predictive Bayesian framework (fortini2019quasi; fortini2025exchangeability; fong2023martingale). In this formulation, the learner’s state $Z_{t}$ is characterised by its predictive distribution over future observations rather than by latent parameters or priors. This distinction is crucial in deep learning, where network parameters serve primarily as computational intermediaries and often lack interpretable statistical meaning. By tracking the evolution of predictions instead of parameter updates, we obtain a formulation of learning that remains empirically interpretable.
A further advantage of this predictive perspective is that it can be validated: predictive statements can be evaluated against realised outcomes, providing an immediate criterion for model adequacy (fong2023martingale). In contrast, probability statements about latent variables, such as priors or posteriors, do not admit direct validation. Framing beliefs in predictive terms also allows for discussion of stationarity with respect to the learner: whether a learner’s sequential forecasts remain self-consistent as learning progresses. This notion of predictive self-consistency provides the conceptual foundation for analysing the dynamics of learning and, ultimately, for our definition of forgetting in § 4.2.
@@ 3.3 Learning as a stochastic process
The stochasticity of the environment induces stochasticity in the learner’s state evolution; as the learner interacts with the environment, the sequence $\{Z_{t}\}$ forms a stochastic process. Since each state $Z_{t}$ defines an induced future distribution $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$, the sequence of induced future distributions over time also forms a stochastic process. Equivalently, the distribution over learner states at time $t$ defines a distribution over distributions of futures, capturing the learner’s evolving uncertainty about possible continuations. This is illustrated in Figure 1.
Takeaway 3.1 The stochastic sequence of observations drives a stochastic evolution of states $\{Z_{t}\}$, which in turn induces a stochastic evolution of induced future distributions $\{q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\}$.
@@
The abstract variables $(X_{t},Y_{t},Z_{t})$ admit natural interpretations across learning settings. In supervised learning, $X_{t}$ denotes (current input, previous target) pairs and $Y_{t}$ the learner’s output in response to the previous input; in RL, $X_{t}$ contains states and rewards while $Y_{t}$ are actions; in generative modelling, $Y_{t}$ directly models $X_{t}$.
The learner state $Z_{t}$ encapsulates all contents of the learner (including parameters, latent variables, buffers, etc.). Distinct modes of operation correspond to different update rules: in training (update rule $u$), both the learner’s belief and auxiliary structures may adapt; in inference ($u^{\prime}$), the learner’s belief remains fixed while transient components continue to evolve.
@@ 4.1 Characterising forgetting
Forgetting is the loss of previously acquired information or capabilities (jagielski2022measuring). To formalise this notion, we first outline desiderata that any valid measure should satisfy, guided by thought experiments (see Appendix C for detailed scenarios).
Desideratum 1.
@@
Forgetting is distinct from task performance; a learner may maintain outdated or incorrect beliefs yet still perform well, or may discard relevant knowledge without an immediate drop in performance. A good measure of forgetting must therefore disentangle information loss from performance fluctuations.
Desideratum 2.
@@
For instance, when a learner revises its predictions in light of new observations, its state changes. This update is not necessarily forgetting, since this state change can occur without any loss of information.
Desideratum 3.
A forgetting measure must capture losses in a learner’s prior capabilities, not just the retention of previously observed data.
Forgetting includes the loss of general capabilities, not just individual observations. For instance, a learner may fail to generalise to unobserved examples that it previously handled correctly. Measures of forgetting that only focus on memorisation overlook this broader notion.
Desideratum 4.
Forgetting should characterise the learner’s loss of information and capabilities resulting from the learner’s adaptations.
Forgetting is a property of the learner, not a property of the environment or dataset in which it operates. For example, a system that never adapts cannot meaningfully forget. Therefore, a learner’s environment may influence the rate or magnitude of forgetting, but it does not define it.
These desiderata provide a principled test bed for developing and evaluating measures of forgetting. The complete set of thought experiments in Appendix C illustrates the need for these desiderata and motivates the formalism developed below.
@@ 4.2 Forgetting
Learning is inherently a process of both absorption and loss. Each new observation provides the learner with new information while simultaneously displacing or overwriting previously learned information (as illustrated in Figure 2). This interdependence is fundamental: every update reshapes what the learner can represent and, consequently, what it may forget.
Figure 2: Information absorption and loss over a single learning update. The figure illustrates the information dynamics occurring within each learning step. At each update, the previous state, $Z_{t-1}$, receives a new observation, $X_{t}$. The learner incorporates information from $X_{t}$ into its state while simultaneously displacing part of the information previously encoded in $Z_{t-1}$. This process results in a new state $Z_{t}$, reflecting the assimilation of new information and the loss of some prior information.
If a learner’s behaviour changes when exposed to data it already expects, it can only be losing information, since no new information has been introduced. The learner’s expectations are defined by its predictive distribution for a given observation. We leverage this intuition to formalise forgetting.
@@
At each time $t$, the state $Z_{t}$ induces a distribution over futures $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$, which characterises both the learner’s current predictive capabilities and its expectations about futures.
We therefore define forgetting in terms of induced futures, rather than the learner’s state or predictions. Different states–whether parametric, non-parametric, or otherwise–may induce identical futures:
@@
where $Z_{t}\neq Z_{t}^{\prime}$. Furthermore, predictions about already-seen observations may shift without indicating that forgetting has occurred. By grounding our definition in induced futures, we ensure that updates are evaluated with respect to the learner’s expected predictions, thereby distinguishing between constructive and destructive adaptation. This satisfies 2 and 3.
@@
Building on this intuition, a learner is unforgetful if updating on targets that are entirely consistent with its own predictions does not change its induced futures. The distribution of futures is expected to remain invariant when no new information is provided to the learner. Notably, suppose a learner’s update function $u(Z_{t-1},X_{t},Y_{t})$ depends on the history $H_{0:t-1}$. In that case, consistency may be violated and forgetting can occur (see § B.3 for a detailed discussion).
Formally, let $q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})$ denote the distribution over futures before updates on targets consistent with the learner’s expectations, and let $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$ denote the distribution after updating on targets consistent with the learner’s expectations. Non-forgetting requires:
$q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{Y_{t},X_{t},Z_{t}}\left[q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\right],$ (7)
where $Y_{t}\sim q_{f}(\cdot\mid Z_{t-1},X_{t-1})$, $X_{t}\sim q_{e}(\cdot\mid H_{0:t-1},Y_{t})$, and $Z_{t}\sim u(\cdot\mid Z_{t-1},X_{t},Y_{t})$.
In these updates, $X_{t}$ is sampled from the same hybrid distribution $q_{e}$ introduced in § 3.2. Updates are performed on these learner-consistent targets to ensure a separation between forgetting and backward transfer. In this formulation, the expectation in (7) is taken over both the stochasticity of the learner’s target generation and the environmental inputs, yielding a description of predictive consistency.
This notion extends naturally to multiple updates. Over $k$ steps, non-forgetting requires that futures induced before any updates remain compatible with those induced after $k$ interactions have occurred. Equivalently, earlier futures must be exactly recoverable by marginalising over all possible $k$-step interaction paths, leading to the generalised consistency condition.
Definition 4.1 (Consistency Condition).
@@
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{X_{t:t^{\prime}},Y_{t:t^{\prime}},Z_{t:t^{\prime}}}\left[q(H^{t+k:\infty}\mid Z_{t^{\prime}},H_{0:t^{\prime}})\right],$ (8)
where $t^{\prime}=t+k-1$, and for $i=t,\ldots,t^{\prime}$ the expectation is taken over
$X_{i}\sim q_{e}(\cdot\mid H_{0:i-1},Y_{i}),\quad Y_{i}\sim q_{f}(\cdot\mid Z_{i-1},X_{i-1}),\quad Z_{i}\sim u(\cdot\mid Z_{i-1},X_{i},Y_{i}).$
@@
In practice, most learning algorithms do forget. To quantify how much the learner is likely to forget at any time $t$, we introduce a concrete measure grounded in our formalism. When the consistency condition is violated, the learner’s updated induced futures diverge from its initial induced futures. Measuring this divergence yields a natural notion of a learner’s propensity to forget. This operationalises our conceptual definition, allowing us to validate the formalism by ensuring that the measure aligns with intuitive expectations about forgetting in practice.
@@
The $k$-step propensity to forget incurred at time $t$ is measured as the divergence
$\Gamma_{k}(t)\coloneqq\mathrm{D}\left(q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\,\|\,q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\right)$ (9)
where $\mathrm{D}(\cdot\|\cdot)$ is a suitable divergence measure.
Takeaway 4.1 Forgetting occurs when the consistency condition is violated; the induced futures after $k$ updates are no longer recoverable from those before the updates.
@@
The perspective of the predictive Bayesian formalism assumes that the learner’s internal state parameterises a coherent predictive distribution over futures: that is, there exists a well-defined mapping
$Z_{t}\mapsto q(H^{t+1:\infty}\mid Z_{t},H_{0:t}),$ (10)
such that all of the learner’s expectations about future interactions arise from a single, self-consistent probabilistic model. This condition ensures that the learner’s predictive state fully captures its beliefs about the future and that updates to these beliefs evolve within a single probability space. When this mapping exists, the induced futures $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$ are well-defined and measurable, allowing the propensity to forget to be meaningfully evaluated. In practice, coherence is typically achieved throughout training, when the learner’s state defines a consistent model of its futures. During certain transitory phases, however, such as buffer reinitialisation, target-network lag, or policy resets, this mapping may temporarily cease to exist. In those intervals, the learner’s state no longer induces a coherent distribution over futures, and the notion of forgetting becomes undefined, not because the formalism fails, but because the learner itself momentarily lacks a predictive model of its behaviour.
More generally, some algorithms may never instantiate such a coherent predictive mapping during their operation. These systems, by construction, do not admit a single probabilistic interpretation of their evolving state and thus fall outside the scope of this formalism. The previously identified conditions of failure (when the learner’s updates do not correspond to transformations within a unified predictive model family) delineate this boundary. In all other cases, the induced futures remain well-defined whenever coherence is achieved, and the forgetting formalism applies without modification.
@@ 5 Empirical analysis
We now examine forgetting empirically. Our theoretical account provides a general conceptualisation of forgetting. To illustrate its utility, we instantiate and study Definition 4.2 across multiple environments and learning algorithms.
@@
Figure 3: Forgetting occurs across all deep learning scenarios. Left: Forgetfulness dynamics of a shallow neural network trained on regression, classification, and generative modelling tasks. The solid line shows the $k$-step forgetfulness (where $k$ varies from 11 to 4040) over the normalised training step. Regression and classification tasks use KL divergence, while the generative task uses the maximum mean discrepancy (MMD). Forgetting dynamics vary throughout training, even without any distribution shift. Right: A class-incremental learning example using a single-layer neural network on a two-moons classification task. We show the $k$-step forgetting profile over four seeds, with the shaded area indicating the spread of $\Gamma_{k}(t)$ over $k$ from 1 to 40. The plot illustrates the abrupt increase in forgetting at the task boundary. See Appendix F for details on the experimental implementation.
5.1 Forgetting profiles in different settings
Across all examined settings, forgetting is non-zero, and both its dynamics and magnitude vary significantly (see Figure 3): this motivates our paper title, “Forgetting is Everywhere”. Absolute $\Gamma_{k}(t)$ values are domain-specific, but the shape and dynamics of the forgetting trajectory convey meaningful information. As shown in Figure 4, even minor changes to the learning algorithm can substantially alter $\Gamma_{k}(t)$.

### v2

@@ 3.2 Predictive distributions
This construction is inspired by predictive Bayesianism (fortini2019quasi; fortini2025exchangeability) and the martingale posterior framework (fong2023martingale). In this setting, an infinite sequence of future observations is simulated, and the resulting statistics form a martingale that can be shown to converge. A central challenge in CL is that learning occurs in parameter space; however, quantities of interest are in data space. Predictive Bayes provides a principled transformation from parameter updates to predictive statements. This allows predictive statements to be validated against realised outcomes and provides an interpretable representation of the learner’s knowledge at any time.
@@ 3.3 Learning as a stochastic process
The stochasticity of the environment induces stochasticity in the learner’s state evolution; throughout the interaction process, the sequence $\{Z_{t}\}$ forms a stochastic process, which in turn defines a stochastic process over predictive distributions $\{q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\}$. Thus, the distribution over learner states at time $t$ defines a distribution over predictive distributions. This is illustrated in Figure 1.
Figure 1: State evolution and predictive distributions. The learner’s internal state $Z_{t}$ evolves through training updates $u$ given $(Z_{t-1},X_{t},Y_{t})$. Each state induces a predictive distribution $q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t})$, which is updated in inference mode $u^{\prime}$. This separation illustrates how each state encodes retained capabilities, and how training vs. introspective updates interact.
@@
The abstract variables $(X_{t},Y_{t},Z_{t})$ admit natural interpretations across learning settings. In supervised learning, $X_{t}$ denotes (current input, previous target) pairs and $Y_{t}$ the learner’s output in response to the previous input; in RL, $X_{t}$ contains states and rewards while $Y_{t}$ are actions; in generative modelling, $Y_{t}$ directly models $X_{t}$. Across these paradigms, the learner state $Z_{t}$ encapsulates all contents of the learner (including parameters, latent variables, buffers, etc.). Distinct modes of operation correspond to different update rules: in training (update rule $u$), both the learner’s belief and auxiliary structures may adapt, whereas an inference update ($u^{\prime}$) leaves its beliefs fixed while transient components continue to evolve.
@@ 4.1 Characterising forgetting
Before formalising forgetting, we first outline desiderata that any valid notion of forgetting should satisfy. These desiderata are motivated by thought experiments, detailed in § C.
Desideratum 4.1.
@@
Forgetting is distinct from measurable success on a task, such as accuracy or cumulative reward. It is a property of the learner itself, independent of the environment. A learner can maintain outdated or incorrect beliefs yet still perform well, or lose relevant knowledge without an immediate change in task outcomes. Conventional metrics, including backward transfer, may fail to capture these underlying changes in knowledge, highlighting the need to disentangle forgetting from task performance.
Desideratum 4.2.
@@
When a learner incorporates new observations, its beliefs (and therefore its state) will change. A change in belief does not necessarily imply anything is forgotten. Therefore, conceptualisations based on changes to beliefs (or parameters) can misidentify information-preserving updates as forgetting.
Desideratum 4.3.
Forgetting should characterise the learner’s loss of prior information and capabilities, not just the retention of previously observed data.
Forgetting encompasses the loss of general capabilities, not only individual observations. For example, a learner may fail to generalise to unobserved examples that it previously handled correctly. Conceptualisations of forgetting that prioritise memorisation overlook this broader notion.
Desideratum 4.4.
Forgetting is a property of the learner, not of the environment in which it operates.
An environment cannot forget; however, it can influence the rate or magnitude of forgetting. These desiderata provide a principled foundation for developing and evaluating conceptualisations of forgetting. The thought experiments in § C justify the desiderata and formalism developed below.
@@ 4.2 Forgetting
Learning is a process of both absorption and loss. Each new observation provides the learner with new information while simultaneously displacing or overwriting previously learned information. Therefore, every update reshapes what the learner can represent and, consequently, what it may forget.
If a learner’s behaviour changes after observing information that it has already incorporated from previous observations, this cannot reflect the acquisition of new information. Therefore, it indicates a loss of prior knowledge. The learner’s expectations are formalised by its predictive distribution; consequently, simulating training updates on samples drawn from this predictive distribution provides a principled perspective on the learner’s likelihood of forgetting at any point in time.
@@
At each time $t$, the state $Z_{t}$ induces a predictive distribution $q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$, which characterises both the learner’s current predictive capabilities and its expectations about futures.
We therefore define forgetting in terms of induced futures, rather than the learner’s state. Different states–whether parametric, non-parametric, or otherwise–may induce identical futures:
@@
where $Z_{t}\neq Z_{t}^{\prime}$. Furthermore, predictions about already-seen observations may shift without indicating that forgetting has occurred. By grounding our definition in induced futures, we ensure that updates are evaluated with respect to the learner’s expected predictions, thereby distinguishing between constructive and destructive adaptation. This satisfies § 4.1 and § 4.1.
@@
Intuitively, a learner is unforgetful if, in expectation, its predictive distribution is invariant to updates on targets that are consistent with its own predictions.
Formally, let $q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ denote the predictive distribution. We define a simulated marginalisation, in which $k$ updates are performed on targets sampled from the learner, as follows.
Definition 4.5 ($k$-Step Simulated Marginalisation).
The $k$-step simulated marginalisation $q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ induced by updating on learner-consistent targets and environmental inputs is given by
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\coloneqq\mathbb{E}_{X_{t:t^{\prime}},Y_{t:t^{\prime}},Z_{t:t^{\prime}}}\left[q(H^{t+k:\infty}\mid Z_{t^{\prime}},H_{0:t^{\prime}})\right],$ (7)
where $t^{\prime}=t+k-1$, and for $i=t,\ldots,t^{\prime}$ the expectation is taken over $X_{i}\sim q_{e}(\cdot\mid Y_{i},Z_{i-1})$, $Y_{i}\sim q_{f}(\cdot\mid Z_{i-1},X_{i-1})$, $Z_{i}\sim u(\cdot\mid Z_{i-1},X_{i},Y_{i})$.
Updates are performed on learner-consistent targets (with $q_{e}$ as introduced in § 3.2) to separate forgetting from backward transfer. In this formulation, the expectation accounts for the stochasticity of target generation, environmental inputs, actions, and state updates.
A learner is non-forgetting if simulated marginalisation results in the same predictive distribution as before the simulated marginalisation $q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$. This yields the following notion of predictive self-consistency.
Definition 4.6 (Self-Consistency Condition).
@@
$q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})=q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1}).$ (8)
Intuitively, this implies that the learner’s predictive distribution must be the same as the distribution recovered from its own simulated marginalisation.
@@
In practice, most learning algorithms do forget. To quantify how much the learner is likely to forget at any time, we introduce a concrete measure grounded in our formalism. When the self-consistency condition (§ 4.2) is violated, the learner’s updated predictive distribution diverges from its initial predictive distribution. Measuring this divergence yields a natural notion of a learner’s propensity to forget, operationalising our conceptual definition.
@@
The $k$-step propensity to forget incurred at time $t$, using a suitable divergence measure $\mathrm{D}(\cdot\|\cdot)$, is given by
$\Gamma_{k}(t)\coloneqq\mathrm{D}\left(q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\,\|\,q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})\right).$ (9)
Takeaway 1.
Forgetting occurs when the self-consistency condition is violated; the predictive distribution after $k$ updates is no longer recoverable from those before the updates.
@@
Our formalism applies whenever the learner’s predictive distribution accurately represents the learner’s state, $Z_{t}\mapsto q(H^{t+1:\infty}\mid Z_{t},H_{0:t})$. Only information used to generate predictions contributes; state components that do not influence predictions (e.g., unused buffer entries) are excluded. Typically, the predictive distribution reflects the state, but this may not be the case during transitory phases such as buffer reinitialisation, target-network lag, or other mechanisms that temporarily decouple the state from predictions. In these intervals, forgetting is undefined, not because the formalism fails, but because the learner temporarily lacks a predictive model of its behaviour. Some algorithms may never produce a predictive mapping and thus fall outside the scope of this formalism. In most cases, however, the predictive distribution is representative of the state.
@@ 5 Empirical analysis
Our theoretical account provides a general conceptualisation of forgetting. To illustrate its utility, we empirically study § 4.2 across multiple environments and learning algorithms.
5.1 Unforgetful learners
Figure 2: Exact Bayesian learners update parameters without forgetting. Axes showing observations $(x,y)$ are shaded white, and parameter axes $(w_{0},w_{1})$ are shaded grey. Top row: The same four observations are presented to a linear regression learner in different orders. Second row: By Theorem 5.1, an exact Bayesian posterior can update its parameters without forgetting. Bottom rows: Two constrained learners – a Gaussian variational posterior with diagonal covariance and a point estimate updated by gradient descent – violate self-consistency, resulting in forgetting. The final column shows the resulting posterior predictive distributions.
Theorem 5.1 (Bayesian Non-Forgetting).
Let $k\in\mathbb{N}$ and for every $t\in\mathbb{N}$ define $t^{\prime}=t+k-1$. Under Assumptions B.2, B.3, B.4, B.1 and B.5 the learner’s posterior-predictive distribution satisfies
$q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})=q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ (10)
when $Z_{t}$ is determined by deterministic Bayesian updates. Thus, $\Gamma_{k}(t)=0$ for all $k,t\in\mathbb{N}$.
We defer the proof of this theorem to § B.3.
Martingale interpretation.
Filtrations formalise the accumulation of information over time in non-forgetting settings. In this perspective, the learner’s posterior-predictive distribution can be seen as a stochastic process $(Q_{t})_{t\geq 0}$ adapted to the state filtration $(\mathcal{G}_{t})_{t\geq 0}$, which encodes all information that has been encoded by the learner up to time $t$. Exact Bayesian updates ensure that $(Q_{t})$ forms a martingale with respect to this filtration $\mathbb{E}[Q_{t+1}\mid\mathcal{G}_{t}]=Q_{t}$ (fong2023martingale). This indicates that, in expectation, updating the learner state on a simulated future observation does not change the predictive distribution. Violations of this property correspond to a drift in the learner’s predictive beliefs, which manifests as forgetting. For a detailed discussion of the martingale perspective on forgetting, see § B.4.
Non-forgetting with parameter updates.
A prevalent conception in continual learning is that updating parameters necessarily causes forgetting (mccloskey1989catastrophic; french1999catastrophic; rusu2016progressive; kirkpatrick2017overcoming; zenke2017continual; aljundi2018memory; masse2018alleviating; li2024spirf; zhao2023does; shenfeld2025rl). However, this conflates changes made to parameters with the loss of the information they represent. Exact Bayesian learners provide a counterexample: although their parameters are updated with every new observation, the resulting posterior is capable of remaining self-consistent. Thus, the learner does not forget despite the changes made to its parameters.
This demonstrates that forgetting is caused by the failure of an update rule to preserve its predictive information (see Figure 2). While many regularisation techniques attempt to mitigate forgetting by penalising parameter changes (kirkpatrick2017overcoming; zenke2017continual; lwf), this only addresses a proxy of the actual issue. In many cases, different parameter configurations admit the same predictive distribution. Ultimately, forgetting reflects a loss of predictive information and can be prevented despite changes made to parameters.
Takeaway 2.
Algorithms may adapt to new observations without forgetting.
5.2 Forgetting in deep learning
Across deep learning settings, forgetting is consistently non-zero, with both its magnitude and dynamics varying substantially (Figure 3). Even in i.i.d. settings, forgetting fluctuates throughout training, reflecting how neural networks continually update and overwrite their knowledge. Although absolute $\Gamma_{k}(t)$ values are domain-specific, the forgetting trajectory reveals how the learner’s beliefs evolve over time. Forgetting is, therefore, functionally meaningful in all tasks, highlighting the importance of a general conceptualisation of forgetting, motivating our paper title, “Forgetting is Everywhere”.
@@ 5.2 Forgetting in deep learning
Figure 3: Forgetting occurs across all deep learning scenarios. Left: Forgetfulness dynamics of a shallow neural network trained on regression, classification, and generative modelling tasks. The solid line shows the $k$-step forgetfulness (where $k$ varies from 11 to 4040) over the normalised training step. Regression and classification tasks use KL divergence, while the generative task uses the maximum mean discrepancy (MMD). Forgetting dynamics vary throughout training, even without any distribution shift. Right: A class-incremental learning example using a single-layer neural network on a two-moons classification task. We show the $k$-step forgetting profile over four seeds, with the shaded area indicating the spread of $\Gamma_{k}(t)$ over $k$ from 1 to 40. The plot illustrates the abrupt increase in forgetting at the task boundary. See § F for details on the experimental implementation.